OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6).

OpenStudy (aroub):

=-6-8/-3+4 =-14/1 =-14

OpenStudy (anonymous):

thanks so much

OpenStudy (aroub):

anytime :D